# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ Abstract
The attention mechanism is the key to the success of transformers in different machine learning tasks. However, the quadratic complexity with respect to the sequence length of the vanilla softmax-based attention mechanism becomes the major bottleneck for the application of long sequence tasks, such as vision tasks. Although various efficient linear attention mechanisms have been proposed, they need to sacrifice performance to achieve high efficiency. What’s more, memory-efficient methods, such as FlashAttention-1-3, still have quadratic computation complexity which can be further improved. In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism to achieve low memory input/output operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is FlashAttention friendly. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism on high-resolution vision tasks without losing performance. Even on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W. The code of ELFATT is available at [https://github.com/Alicewithrabbit/ELFATT].
Index Terms:
Efficient attention mechanism, linear approximation, vision transformer.
I Introduction
@@
Transformers have achieved great success in large language models (ChatGPT [1] and Llama [2]) and large vision models (SAM [3] and Sora [4]). The core technique of the transformer, the vanilla softmax-based attention mechanism, is capable of capturing the relationship between any two tokens [5]. In complex tasks, the sequence length becomes longer and longer, usually much longer than the embedding dimension. The softmax operation after the multiplication of query and key matrices makes the vanilla softmax-based attention mechanism quadratic computational complexity with respect to the sequence length. It has become a main obstacle to computational efficiency.
@@
Acceleration methods to speed up the attention computation can be classified into two categories: (1) memory-efficient methods; (2) computation-efficient methods. Memory-efficient methods focus on optimizing memory input/output (I/O) operations to achieve almost linear complexity [6, 7, 8, 9]. FlashAttention [6, 7, 8] and FlashSigmoid [9] are representatives of memory-efficient methods. Computation-efficient methods focus on minimizing the computation bound of the attention computation by using linear approximations based on different kernel methods [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], low-rank decomposition [20, 21, 15, 5], and sparse computation [22, 23, 24, 25, 26, 27, 28, 29, 30]. However, memory-efficient methods still have quadratic computational complexity, and computation-efficient methods usually have lower performance compared to the vanilla softmax-based attention mechanism.
@@
In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to the vanilla softmax-based attention mechanism. The core idea of ELFATT is the combination of sparse computation with a linear approximation. Each ELFATT block has two parallel attention heads. One head is used to compute sparse blockify attention to introduce inductive biases, and the other one head is used to compute global linear attention to capture long-range dependencies. Both heads have almost linear complexity, and the sparse blockify attention process can be further speeded up by using the FlashAttention [6, 7, 8] mechanisms to reduce memory I/O operations. ELFATT is evaluated on different vision tasks: image classification, semantic segmentation, and object detection. What’s more, ELFATT is also evaluated on edge GPUs. Compared to state-of-the-art memory-efficient acceleration methods and computation-efficient acceleration methods, ELFATT inherits advantages of both two kinds of methods: noninferior performance compared to the vanilla softmax-based attention mechanism, low memory I/O operations, and linear computational complexity. In summary, this paper has the following contributions.
@@
    A novel efficient linear fast attention (ELFATT) mechanism is proposed that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to the vanilla softmax-based attention mechanism.
@@
    The relationship between the proposed ELFATT mechanism and the vanilla softmax-based attention mechanism is analyzed and given.
@@
    An upper bound is given for the use of the proposed ELFATT mechanism to approximate the vanilla softmax-based attention mechanism.
@@
    ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism without using FlashAttention-2 and 2-3x speedups over the vanilla softmax-based attention mechanism using FlashAttention-2 on high-resolution vision tasks.
@@
    ELFATT offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W on edge GPUs.
II Related Work
@@
Since the vanilla softmax-based attention mechanism has quadratic complexity in both time and memory, depending on whether the focus is on optimizing memory complexity or time complexity, attention acceleration methods can be categorized into the following two types: (A) memory-efficient attention acceleration methods; (B) computation-efficient attention acceleration methods.
II-A Memory-Efficient Attention Acceleration Methods
@@
Memory-efficient attention acceleration methods focus on optimizing quadratic memory complexity to almost linear complexity. One of the most representative methods is FlashAttention [6], which minimizes memory I/O access for the softmax computation of the product of query and key matrices to improve the utilization rate of GPUs to achieve almost linear memory complexity [6, 5]. FlashAttention-2 further reduces the number of floating point operations (FLOPs) of non-matrix multiplication, parallelizes both forward and backward processes according to the sequence length, and reduces the I/O access of shared memory [7]. FlashAttention-3 is proposed to fully utilize the performance of new Hopper GPUs by introducing warp-specialization, interleave block-wise matrix multiplication and softmax operations, and the support of FP8 precision [8]. FlashSigmoid investigates the feasibility of using the sigmoid function to replace the softmax function in attention computation, proposes a new regularity method for sigmoid attention to stabilize training, and introduces a memory-efficient version based on FlashAttention-2[9]. However, these memory-efficient attention acceleration methods are usually hard-aware and can not support all kinds of GPUs and their computation complexity is still quadratic.
II-B Computation-Efficient Attention Acceleration Methods
@@
Different to hard-aware memory-efficient attention acceleration methods, computation-efficient attention acceleration methods focus on optimizing computational complexity by linear approximations and usually can be categorized into three classes: (a) kernel methods; (b) low-rank decomposition based methods; (c) sparse methods.
@@
Kernel methods usually change the order of nonlinear normalization and multiplication of the query matrix and key matrix to achieve linear computational complexity. FLatten [10] introduces a cubic rectified linear unit (ReLU) [31] based feature map and uses it to perform nonlinear normalization on the query matrix and key matrix, respectively, before the attention computation. The linear transformer [13] introduces a feature map based on the exponential linear unit (elu) [32] activation function. cosFormer [17] introduces a cos-based nonlinear feature map to obtain a linear approximation of the vanilla softmax-based attention mechanism. Efficient attention [18] performs softmax normalization on the query matrix and the key matrix, respectively, to achieve linear complexity [18]. However, this linearization introduces noise from two aspects: (1) Separated softmax normalization will reduce the similarity of elements in the query matrix and key matrix that are both negative at the same position; (2) Large negative values are suppressed to small positive values. The above two types of noise result in decreased discrimination between tokens and incorrect concentration of attention. Agent [19] introduces a small agent matrix, which is obtained by performing the pooling operation on the query matrix. This agent matrix is used as the auxiliary key matrix to be multiplied with the query matrix to reduce its dimenison before the softmax normalization. Similarly, it is also used to reduce the dimension of the key matrix before its softmax normalization. This agent matrix served as a bridge between two independent softmax normalization processes to reduce noise.
Low-rank decomposition based methods usually take the softmax normalization of the product of query and key matrices as a whole for decomposition. Nyströmformer [21] introduces Nyström approximation to perform low-rank decomposition of the softmax normalization of the product of query and key matrices. For calculating an attention score matrix, it needs to perform an iterative inverse approximation and operate softmax normalization three times (each softmax normalization still involves a product of two matrices), which makes its acceleration effect not significant. Skeleton decomposition-based self-attention uses a simplified inverse approximation method based on the permuted diagonal matrix [20]. Interactive multihead self-attention (iMHSA) introduces several linear layers to perform head interactions instead of performing an iterative inverse approximation [33]. CUR decomposition based self-attention (CURSA) [5] introduces the CUR decomposition to replace the Nyström approximation and reduces the number of matrix multiplication.
@@
Sparse methods usually separate the original sequences into smaller blocks using sliding windows and perform attention computation within these blocks to reduce complexity. Each token in a sequence is only affected by several tokens, not all tokens; hence, this kind of sparse methods are also called local attention mechanism. Swin [22] and Longformer [26] are pioneers in introducing sliding windows into attention computation for vision tasks and language tasks, respectively. To address the information loss introduced by the local windows, Swin introduces a shifted window mechanism to capture cross-block information, while Longformer selects some tokens as global tokens, which have effects on all tokens, to compensate for the global information loss. NesT [34] further simplifies the shifted window mechanism through simple spatial operations. CSWin [23] introduces a cross-shaped window mechanism for 2D sequences. The cross-shaped window mechanism separates a 2D sequence into horizontal and vertical stripes, respectively, and performs parallel attention computation within these stripes in horizontal and vertical directions, which can simultaneously obtain local inductive biases and cross-block (global) information. In addition to using the deterministic global token selection mechanism of Longformer, Big Bird [29] also randomly selects some global tokens to enhance performance. Similarly, the sparse transformer [27] proposed by OpenAI also introduces several fixed-step tokens, which is similar to dilated convolution [35] in convolutional neural networks (CNNs) to capture long-range information. Reformer [30] finds the nearest neighbors for each token to calculate local attention scores using the locality-sensitive-hashing (LSH) algorithm.
III Methods
III-A Vanilla Softmax-Based Attention Mechanism
@@
For clarity, we assume that all vectors appear in this paper are row vectors. For any two same sized tokens $\textit{{x}}\in\mathbb{R}^{c}$ and $\textit{{y}}\in\mathbb{R}^{c}$, their attention similarity score can be calculated as follows,
$a={\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right),$ (1)
@@
where ${\rm exp}\left(\cdot\right)$ is an element-wise exponential function and is conducted after the inner product of two tokens. For given two $m$-tuples $\{\textit{{x}}_{1},\dots,\textit{{x}}_{m}\}$ and $\{\textit{{y}}_{1},\dots,\textit{{y}}_{m}\}$ with $\textit{{x}}_{i},\textit{{y}}_{i}\in\mathbb{R}^{c}$ and $i=1,2,\dots,m$, we need to calculate $m^{2}$ pairs of attention similarity scores. Currently, a long sequence length $m$ is preferred in large models and $m\gg c$, therefore, the computational complexity of the vanilla softmax-based attention mechanism is quadratic with respect to the sequence length $m$.
III-B General Attention Mechanism
Eq. (1) can be rewritten as a more general form [13] as follows,
$a={\rm sim}\left(\textit{{x}},\textit{{y}}\right),$ (2)

### v4

@@ Abstract.
The attention mechanism is the key to the success of transformers in different machine learning tasks. However, the quadratic complexity with respect to the sequence length of the vanilla softmax-based attention mechanism becomes the major bottleneck for the application of long sequence tasks, such as vision tasks. Although various efficient linear attention mechanisms have been proposed, they need to sacrifice performance to achieve high efficiency. What’s more, memory-efficient methods, such as FlashAttention-1-3, still have quadratic computation complexity which can be further improved. In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism to achieve low memory input/output operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is FlashAttention friendly. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism on high-resolution vision tasks without losing performance. Even in some non-vision tasks of long-range arena, ELFATT still achieves leading performance and offers 1.2-2.3x speedups over FlashAttention-2. Even on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W. Furthermore, ELFATT can be used to enhance and accelerate diffusion tasks directly without training.
Efficient Attention Mechanism; Linear Approximation; Vision Transformer
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754825isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Machine learningccs: Theory of computation Design and analysis of algorithmsccs: Computing methodologies Neural networksccs: Computing methodologies Computer vision
1. Introduction
@@
Transformers have achieved great success in large language models (ChatGPT (Achiam et al., 2023) and Llama (Grattafiori et al., 2024)) and large vision models (SAM (Kirillov et al., 2023) and Sora (Brooks et al., 2024)). The core technique of the transformer, the vanilla softmax-based attention (VaniATT) mechanism, is capable of capturing the relationship between any two tokens (Wu et al., 2024). In complex tasks, the sequence length becomes longer and longer, usually much longer than the embedding dimension. The softmax operation after the multiplication of query and key matrices makes VaniATT quadratic computational complexity with respect to the sequence length. It has become a main obstacle to computational efficiency.
@@
Acceleration methods to speed up the attention computation can be classified into two categories: (1) memory-efficient methods; (2) computation-efficient methods. Memory-efficient methods focus on optimizing memory input/output (I/O) operations to achieve almost linear complexity (Dao et al., 2022; Dao, 2024; Shah et al., 2024; Ramapuram et al., 2025). FlashAttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and FlashSigmoid (Ramapuram et al., 2025) are representatives of memory-efficient methods. Computation-efficient methods focus on minimizing the computation bound of the attention computation by using linear approximations based on different kernel methods (Han et al., 2023; Choromanski et al., 2021; Wortsman et al., 2023; Katharopoulos et al., 2020; Bolya et al., 2023; Lu et al., 2021; Peng et al., 2021; Qin et al., 2022; Shen et al., 2021; Han et al., 2025; Ran et al., 2025) or low-rank decomposition (Han et al., 2022; Xiong et al., 2021; Lu et al., 2021; Wu et al., 2024), and sparse computation (Liu et al., 2021; Dong et al., 2022; Wang et al., 2020a; Zhao et al., 2019; Beltagy et al., 2020; Child et al., 2019; Tay et al., 2020; Zaheer et al., 2020; Kitaev et al., 2020). However, memory-efficient methods still have quadratic computational complexity. And computation-efficient methods usually have lower performance compared to VaniATT (Liu et al., 2023).
@@
In this paper, we propose a novel efficient linear fast attention (ELFATT [^1] The ELFATT: Efficient Linear Fast Attention for Vision Transformers methodology disclosed in this work is protected under intellectual property rights, including the following patent applications and granted patents: U.S. Non-Provisional Patent Application (No. 19/011,779 filed on 7 January 2025), Hong Kong Short-Term Patent (No. HK30118021A2 granted on 30 May 2025), and Chinese Invention Patent Application (No. 202510108088.7 filed on 23 January 2025). Commercial inquiries should be directed to kt@innocimda.com.) mechanism that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to VaniATT. The core idea of ELFATT is the combination of sparse computation with a linear approximation. Each ELFATT block has two parallel attention heads. One head is used to compute sparse blockify attention to introduce inductive biases, and the other one head is used to compute global linear attention to capture long-range dependencies. Both heads have almost linear complexity, and the sparse blockify attention head can be further speeded up by using the FlashAttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) mechanisms to reduce memory I/O operations. ELFATT is evaluated on different vision tasks: image classification, semantic segmentation, object detection, and diffusion. Furthermore, ELFATT is also evaluated on edge GPUs and in the long-range arena. Compared to state-of-the-art memory-efficient acceleration methods and computation-efficient acceleration methods, ELFATT inherits advantages of both two kinds of methods: noninferior performance compared to VaniATT, low memory I/O operations, and linear computational complexity. In summary, this paper has the following contributions.
@@
    A novel efficient linear fast attention (ELFATT) mechanism is proposed that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to VaniATT.
@@
    ELFATT is the first sparse blockify attention + global linear attention design for attention acceleration which is significantly different from the existing hierarchical sparse attention mechanisms.
@@
    The relationship between ELFATT and VaniATT is analyzed and given. An upper bound is given for the use of ELFATT to approximate VaniATT. To the best of our knowledge, we are the first to provide a theoretical bound analysis to sparse attention.
@@
    ELFATT offers 4-7x speedups over VaniATT without using FlashAttention-2 and 2-3x speedups over VaniATT using FlashAttention-2 on high-resolution vision tasks.
@@
    ELFATT can be used to accelerate non-vision tasks with 1K-4K sequence lengths and offers 1.2-2.3x speedups over FlashAttention-2.
  • (vi)
@@
    ELFATT offers 1.6x to 2.0x speedups over state-of-the-art attention mechanisms in various power modes from 5W to 60W on edge GPUs.
  • (vii)
    ELFATT can be used to enhance and accelerate diffusion tasks directly without training.
2. Related Work
@@
Since VaniATT has quadratic complexity in both time and memory, depending on whether the focus is on optimizing memory complexity or time complexity, attention acceleration methods can be categorized into the following two types: (A) memory-efficient attention acceleration methods; (B) computation-efficient attention acceleration methods.
2.1. Memory-Efficient Attention Acceleration Methods
@@
Memory-efficient attention acceleration methods focus on optimizing quadratic memory complexity to almost linear complexity. One of the most representative methods is FlashAttention (Dao et al., 2022), which minimizes memory I/O access for the softmax computation of the product of query and key matrices to improve the utilization rate of GPUs to achieve almost linear memory complexity (Dao et al., 2022; Wu et al., 2024). FlashAttention-2 further reduces the number of floating point operations (FLOPs) of non-matrix multiplication, parallelizes both forward and backward processes according to the sequence length, and reduces the I/O access of shared memory (Dao, 2024). FlashAttention-3 is proposed to fully utilize the performance of new Hopper GPUs by introducing warp-specialization, interleave block-wise matrix multiplication and softmax operations, and the support of FP8 precision (Shah et al., 2024). FlashSigmoid investigates the feasibility of using the sigmoid function to replace the softmax function in attention computation, proposes a new regularity method for sigmoid attention to stabilize training, and introduces a memory-efficient version based on FlashAttention-2 (Ramapuram et al., 2025). However, these memory-efficient attention acceleration methods are usually hard aware and can not support all kinds of GPUs and their computation complexity is still quadratic.
2.2. Computation-Efficient Attention Acceleration Methods
@@
Different to hard aware memory-efficient attention acceleration methods, computation-efficient attention acceleration methods focus on optimizing computational complexity by linear approximations and usually can be categorized into two classes: (a) kernel methods; (b) sparse methods.
@@
Kernel methods usually change the order of nonlinear normalization and multiplication of the query matrix and key matrix to achieve linear computational complexity. FLatten (Han et al., 2023) introduces a cubic rectified linear unit (ReLU) (Nair and Hinton, 2010) based feature map and uses it to perform nonlinear normalization on the query matrix and key matrix, respectively, before the attention computation. The linear transformer (Katharopoulos et al., 2020) introduces a feature map based on the exponential linear unit (ELU) (Clevert, 2015) activation function. cosFormer (Qin et al., 2022) introduces a cos-based nonlinear feature map to obtain a linear approximation of VaniATT. Efficient attention (Shen et al., 2021) performs softmax normalization on the query matrix and the key matrix, respectively, to achieve linear complexity. However, this linearization introduces noise from two aspects: (1) Separated softmax normalization will reduce the similarity of elements in the query matrix and key matrix that are both negative at the same position; (2) Large negative values are suppressed to small positive values. The above two types of noise result in decreased discrimination between tokens and incorrect concentration of attention. Agent (Han et al., 2025) introduces a small agent matrix, which is obtained by performing the pooling operation on the query matrix. This agent matrix is used as the auxiliary key matrix to be multiplied with the query matrix to reduce its dimenison before the softmax normalization. Similarly, it is also used to reduce the dimension of the key matrix before its softmax normalization. This agent matrix serves as a bridge between two independent softmax normalization processes to reduce noise. Another type of kernel methods introduce the low-rank decomposition (Xu et al., 2024b, a; Xu and Wei, 2024). They usually take the softmax normalization of the product of query and key matrices as a whole for decomposition to find appropriate feature maps to reduce complexity (Wu et al., 2024). Nyströmformer (Xiong et al., 2021) introduces Nyström approximation to perform low-rank decomposition of the softmax normalization of the product of query and key matrices. For calculating an attention score matrix, it needs to perform an iterative inverse approximation and operate softmax normalization three times (each softmax normalization still involves a product of two matrices), which makes its acceleration effect not significant. Skeleton decomposition-based self-attention uses a simplified inverse approximation method based on the permuted diagonal matrix (Han et al., 2022). Interactive multihead self-attention (iMHSA) introduces several linear layers to perform head interactions instead of performing an iterative inverse approximation (Kang et al., 2024). CUR decomposition based self-attention (CURSA) (Wu et al., 2024) introduces the CUR decomposition to replace the Nyström approximation and reduces the number of matrix multiplication. Mamba (Gu and Dao, 2024) and its variants (Dao and Gu, 2024; Liu et al., 2024; Zhu et al., 2024) are also a kind of kernelized attention mechanism.
@@
Sparse methods usually separate the original sequences into smaller blocks using sliding windows and perform attention computation within these blocks to reduce complexity. Each token in a sequence is only affected by several tokens, not all tokens; hence, this kind of sparse method is also called the local attention mechanism. Swin (Liu et al., 2021) and Longformer (Beltagy et al., 2020) are pioneers in introducing sliding windows into attention computation for vision tasks and language tasks, respectively. To address the information loss introduced by the local windows, Swin introduces a shifted window mechanism to capture cross-block information, while Longformer selects some tokens as global tokens, which have effects on all tokens, to compensate for the global information loss. NesT (Zhang et al., 2022) further simplifies the shifted window mechanism through simple spatial operations. CSWin (Dong et al., 2022) introduces a cross-shaped window mechanism for 2D sequences. The cross-shaped window mechanism separates a 2D sequence into horizontal and vertical stripes, respectively, and performs parallel attention computation within these stripes in horizontal and vertical directions, which can simultaneously obtain local inductive biases and cross-block (global) information. In addition to using the deterministic global token selection mechanism of Longformer, Big Bird (Zaheer et al., 2020) also randomly selects some global tokens to enhance performance. Similarly, the sparse transformer (Child et al., 2019) proposed by OpenAI also introduces several fixed-step tokens, which is similar to dilated convolution (Yu, 2015) in convolutional neural networks (CNNs) to capture long-range information. Reformer (Kitaev et al., 2020) finds the nearest neighbors for each token to calculate local attention scores using the locality-sensitive-hashing (LSH) algorithm.
3. Methods
3.1. Vanilla Softmax-Based Attention Mechanism
@@
For clarity, we assume that all vectors appear in this paper are row vectors. Scaling and normalization are omitted in this paper for convenience. For any two same sized tokens $\textit{{x}}\in\mathbb{R}^{c}$ and $\textit{{y}}\in\mathbb{R}^{c}$, their attention similarity score can be calculated as follows,
(1) $a={\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right),$
@@
where ${\rm exp}\left(\cdot\right)$ is an element-wise exponential function and is conducted after the inner product of two tokens. For given two $m$-tuples $\{\textit{{x}}_{1},\dots,\textit{{x}}_{m}\}$ and $\{\textit{{y}}_{1},\dots,\textit{{y}}_{m}\}$ with $\textit{{x}}_{i},\textit{{y}}_{i}\in\mathbb{R}^{c}$ and $i=1,2,\dots,m$, we need to calculate $m^{2}$ pairs of attention similarity scores. Currently, a long sequence length $m$ is preferred in large models and $m\gg c$, therefore, the computational complexity of VaniATT is quadratic with respect to the sequence length $m$.
3.2. General Attention Mechanism
Eq. (1) can be rewritten as a more general form (Katharopoulos et al., 2020) as follows,
(2) $a={\rm sim}\left(\textit{{x}},\textit{{y}}\right),$